[PRad] RC weekly report

gasparan at jlab.org gasparan at jlab.org
Sun May 15 15:20:11 EDT 2016

Ran Coordinator Report for the week May 9 to May 15

1)	Updates for the Tagger electronics are completed. All T- and E-counters
(except few) are connected and ready for the experiment.
2)	Gas hydrogen at 15 degree was tested was tested for the first time in
the Target. There are some issues on cell to chamber differential
pressures. Target group will continue tests from Monday.
3)	GEM cabling is completed. There are some issues with the GEM DAQ and
the el. magnetic noise level on the signal. Work is in progress on these
4)	 HyCal cooling down is completed, it is running at 16 degrees and ready
for the experiment.
5)	Full DAQ integration completed yesterday (Saturday) evening. Some more
tests are needed.
6)	Planned tests on the on-line programs still waiting for a fully
functioning DAQ.
7)	No beam from Friday afternoon.


More information about the PRad mailing list